OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

So, let x be the number of humans and y be the number of horses. Then the number of heads would be \[x + y = 74\] and the number of legs would be \[2x + 4y = 196\] multiplying the first equation by two and subtracting it from the second equation gives \[2y = 196-144 = 52\] which means that \[y = 26, x = 74-26 = 48\] so there were 48 humans and 26 horses.

OpenStudy (anonymous):

i found the answer on my teachers desk .. its 24 horses and 50 humans. Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 So, the solution is 24 horses and 50 humans.

OpenStudy (anonymous):

oh, oops, that should have been 148 not 144 :) my fault.

OpenStudy (anonymous):

its ok thanks any wayss